(in millions of dollars)

Consumption

20000

Investment

2500

Depreciation

2000

Corporate Profits

2800

Exports

2000

Surpluses of Government Enterprises

2000

Compensation of Employees

9300

Proprietor's Income

2000

Net Interest

30

Government Purchases

800

Direct Taxes

900

Net Indirect Taxes

600

Imports

90

Retained Earnings of Corporations

600

Receipt of Factor Income from Rest of the World

900

Payment of Factor Income to Rest of the World

400

  1. Calculate GDP by income approach.
  2. Calculate National Income (NNP) by expenditure approach.
  3. Calculate Disposable Personal Income.
Answer #1

GDP at MP by income method= wages+ rent+ interest+ Corporate profit+ Proprietor income+compensation of employees +Depreciation+ indirect business tax

=0+0+30+2800+2000+9300+2000+600

=16,730

2.

NNP at FC by expenditure method=C+I+G+X-M-depreciation+ receipt of factor income from Rest of the world – payment of factor income to rest of the world.

=20,000+2500+800+2000-90-2000+900-400

=$23,710

c.

Personal Income=National income+ transfer payments- corporate retained earning- corporate taxes- social security taxes

=23,710+0-600-0-0

=$23,110

Personal Disposable income= Personal income- personal taxes

=23,110-900

=22,210